40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 1 13:56:52 2009, Alberto, Update, PSL, Laser Power after fixing the laser chiller MOPAtrend.jpg
    Reply  Mon Jun 1 15:01:31 2009, rana, Update, PSL, Laser Power after fixing the laser chiller: more traces Picture_7.png
       Reply  Mon Jun 1 19:19:26 2009, rana, Update, PSL, 1000 days of hour-trend u.png
Message ID: 1636     Entry time: Mon Jun 1 13:56:52 2009     Reply to this: 1639
Author: Alberto 
Type: Update 
Category: PSL 
Subject: Laser Power after fixing the laser chiller 

The laser power seems to have become more stable after fixing the laser chiller. The power is lower than it used to be (MOPA amplitude 2.5 versus 2.7) but, as shown in the attchement, it became more steady.

Attachment 1: MOPAtrend.jpg  39 kB  | Hide | Hide all
MOPAtrend.jpg
ELOG V3.1.3-